OpenStudy (anonymous):

Graph with a line going through points four comma 0 and four comma 2. Select the equation of the line that passes through the point (−2, 3) and is perpendicular to the line on the graph. y = 0 y = −2 y = 3 y = 3x My Choice is B Plz Help :)

OpenStudy (trojanpoem):

 Equation of line: (Slope = m = (delta y)/4-4 = 1/ 0 undefined) x = 4 the product of both lines slopes = -1 (if perpendicular) m * p = -1 p = -1/m = -1/1/0 = 0 b = 3 y = mx +b y = 0 + b y = b at (-2,3) b =3 y = 3